OpenStudy (anonymous):

81 to the power of 1/2

OpenStudy (anonymous):

now again ^1/2 means square root hence \[\sqrt{81} = 9\]

OpenStudy (anonymous):

9

OpenStudy (anonymous):

or 81^1/2=9^2^1/2=9

OpenStudy (angela210793):